The length of a rectangle is twice it's width. If the perimeter of the rectangle is 54ft, find it's area.
Answer by ewatrrr(24785)

 
Hi,

p = 2L + 2w = 54ft
0r L + w = 27 L = 2w
2w + w = 27
3w = 27
w = 9ft and L = 18ft
A = Lw = (18ft)(9ft) = 162ft^2